You are a financial planner and your new client, Ying Li, needs your advice. Ying has...

  1. You are a financial planner and your new client, Ying Li, needs your advice. Ying has been working for a few years but has no savings. Given her new promotion and salary increase, she wants to start saving. Ying wants to save $32,000 a year for the next 30 years. She wants to begin withdrawing $180,000 per year from her savings account starting a year after retirement. In addition, to celebrate her 5th year of her retirement, she wants to take a year-long trip with her friends around the world, which is expected to cost $60,000. Ying wants to leave on her own for the next 20 years. After 20 years, her son has agreed to take care of all her expenses, but she wants to give him at least $500,000 to help with the cost. Assuming she can earn 6% on her savings, can she afford this retirement plan? Support your conclusion with appropriate figures (Hint: this problem has several parts with each one focusing of a different sets of withdrawals).
Answer #1

Savings at retirement: PMT = 32,000; N = 30; rate = 6%, solve for FV.

FV = 2,529,861.96

Withdrawals:

Present Value (PV) at retirement for the annual withdrawals of 180,000: PMT = 180,000; N = 20; rate = 6%, solve for PV.

PV1 = 2,064,585.82

PV at retirement of the trip in year 5 = 60,000/(1+6%)^5 = 44,835.49 (PV2)

PV at retirement of the lump sum to be given to her son at the end of year 20 = 500,000/(1+6%)^20 = 155,902.36 (PV3)

Total requirement at retirement = PV1 + PV2 + PV3

= 2,064,585.82 + 44,835.49 + 155,902.36 = 2,265,323.67

Her savings at retirement will be greater than her requirements so she can afford this retirement plan.